Circular No. 6628

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997bt IN ANONYMOUS GALAXY D. Hardin reports her discovery, as part of the EROS team’s system-

atic CCD search with the European Southern Observatory’s Marly 1-m telescope, of a supernova (V ∼ 19.5) on Mar. 31, located 0′′.3 east and 0′′.3 south of the center of a galaxy at α = 11^{h}06^{m}15^{s}.1, δ = −11^{o}36′25′′ (equinox 2000.0). No object was present at the position of the supernova on a CCD frame taken on Feb. 13. The host galaxy, number 2915 in the LCRS Survey, has R_{c} = 16.17 and redshift 0.06. A spectrum obtained by S. Benetti on Apr. 15 at the Danish 1.54-m reflector (+ DFOSC) at La Silla shows features consistent with those of a type-II supernova a few weeks after maximum.

GRB 970402 H. Pedersen, Copenhagen University Observatory; A. Castro-Tirado

and J. Gorosabel, Laboratorio de Astrofísica Espacial y Física Fundamental, INTA, Madrid; L. Metcalfe, European Space Agency, VILSPA, Madrid; L. Nicastro and E. Palazzi, ITESRE, CNR, Bologna; F. Frontera, Università di Ferrará; L. Piro, M. Feroci, and E. Costa, Istituto di Astrofisica Spaziale, Rome; J. Heise, Space Research Organization of the Netherlands, Utrecht; M. V. Alonso, Observatorio de Córdoba; M. Boer, Centre d’Etude Spatiale des Rayonnements, Toulouse; and G. Mathys, M. Scodeggio, J. Brewer, S. Benetti, and F. Comeron, European Southern Observatory (ESO), write: “We have obtained deep images using the ESO New Technology Telescope with the EMMI (9′ × 9′ field, on Apr. 4, 5, and 6) and SUSI (2′.2 × 2′.2 field, on Apr. 7 and 11) instruments. The EMMI data are centered on the 3′-error GRB position (IAUC 6610), while the SUSI data are centered the 50′′-error x-ray position (IAUC 6617). No variable object was found within the area covered by SUSI; we estimate that the limits are V = 22.5 and I = 21.7 for the detection of an emerging or disappearing object. In the area covered by EMMI only (thus well removed from the x-ray source), we found one variable, starlike object, at α = 14^{h}49^{m}43^{s}.6, δ = −69^{o}18′55′′ (equinox 2000.0); its magnitudes were Apr. 4.2, I = 17.25, V = 18.33, B = 19.32, R = 17.86; Apr. 5.3, I = 17.20, B = 19.30; Apr. 6.4, I = 17.76. We note that BL Cir (α = 14^{h}49^{m}56^{s}.7, δ = −69^{o}20′50′′, equinox 2000.0) is close to the x-ray position. On Apr. 4, we found V = 19.6, B = 21.6 for this semi-regular variable; a low-dispersion spectrum obtained on Apr. 5 using the 1.54-m Danish telescope (+ DFOSC) indicates that it is a deeply reddened star of spectral type late-G to early-K.”

1997 April 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green